[Brahms-l] Shift report 20040203 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Tue Feb 03 2004 - 08:00:41 EST
------------------------------------------------------------
Date: Tue Feb  3 08:00:41 2004, Shift: 00:00-08:00
Supervisor: BB, with: 
Log book#15 pages 48 - 48

Shift Summary:
------------------------------------------------------------
1. Beam for 4.5 hours
2. All runs with FS:3 degrees 3450 B, MRS at 40 dgrees 2255 A
   Run 10044:                300k trigger 6
   Run 10045:142k trigger 3, 300k trigger 6
   Run 10046:160k trigger 3, 350k trigger 6
   Run 10047:126k trigger 3, 269k trigger 6
3. Magnets were well behaved. T5-2-all planes and T5-3-all planes showed many missing wires for runs starting at least as far back as 10033. Bob Scheetz suggested recycling the remote power to T5. This worked beginning with run 10047. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 3 08:01:06 2004

This archive was generated by hypermail 2.1.8 : Tue Feb 03 2004 - 08:01:21 EST